OpenStudy (anonymous):

What is the solution to this sytem of equations? 4x+3y=12 7x+5y=26 1. (18,20) 2. (18,-20) 3. (-20,18) 4. (20,18)

OpenStudy (solomonzelman):

I would multiply the first equation times 5, and second equation times -3. Then, I would add the equations and this will help me eliminate one of the variables.

OpenStudy (123qwet12):

I got 18, -20

OpenStudy (solomonzelman):

times 3 \(\color{#000000 }{ \displaystyle 4x+3x=12\quad \Longrightarrow \Longrightarrow \quad (4x\color{red}{\times 3}) +(3x\color{red}{\times 3})=(12\color{red}{\times 3}) }\)

OpenStudy (anonymous):

Yes I got it, thank you guys!

OpenStudy (123qwet12):

For future reference you'll need to use the linear combination method to solve some systems